Reports 1-1 of 1 Clear search Modify search
AdV-PSL (Commissioning of SS)
flaminio - 13:06 Wednesday 07 April 2021 (51342) Print this report
ML without thermal correction
Since the IMC loop was restarted yesterday, the ML thermal correction provided by BsX is constant (see attached plot). The PZT is doing the job of keeping the master laser locked to the IMC. Indeed now the PZT correction is not varying around 0 as it does when the thermal correction is working fine. It seems that something is not working properly in the BsX.
Images attached to this report
Comments to this report:
chiummo - 14:48 Wednesday 07 April 2021 (51343) Print this report

I had forgotten the ML TH correction open, I restored it few minutes ago.

Images attached to this comment
Search Help
×

Warning

×